### v1

# KeyB2: Selecting Key Blocks is Also Important for Long Document Ranking with Large Language Models

Abstract.
The rapid development of large language models (LLMs) like Llama has significantly advanced information retrieval (IR) systems. However, using LLMs for long documents, as in RankLLaMA, remains challenging due to computational complexity, especially concerning input token length. Furthermore, the internal mechanisms of LLMs during ranking are still not fully understood. In this paper, we first explore the internal workings of LLMs during relevance judgement and identify that specific attention heads play a crucial role in aligning relevant tokens. This observation inspires us to revisit the block pre-ranking strategy used in KeyB, which remains state-of-the-art (SOTA) on the TREC 2019 DL document ranking dataset. Building on these insights, we develop KeyB2, an advanced long document IR approach that integrates block pre-ranking with the performance of LLMs. KeyB2 efficiently identifies and processes the most relevant blocks, reducing computational costs and improving ranking effectiveness. Additionally, we introduce a new bi-encoder block matching strategy for KeyB2. Comprehensive experiments on long-document datasets, including TREC 2019 DL, Robust04, and MLDR-zh, show that KeyB2 outperforms baselines like RankLLaMA and KeyB by reducing reranking time and GPU memory usage while enhancing retrieval performance, achieving new SOTA results on TREC 2019 DL with higher NDCG@10 and MAP scores.
Long-document neural information retrieval, large language models
copyright: nonejournalyear: 202xpublicationmonth: 1doi: noneccs: Information systems Information retrievalccs: Information systems Retrieval models and rankingccs: Information systems Language modelsccs: Information systems Document representation
1. Introduction
The rapid development of large language models (LLMs) such as GPT and Llama has transformed diverse natural language processing (NLP) tasks, including information retrieval (IR), text generation, and comprehension. These models have demonstrated remarkable capabilities in capturing deep semantic relationships within text, leading to significant advancements in IR systems. Cross-encoder models, such as RankLLaMA (Ma et al., 2024), have been particularly effective at evaluating the relevance of queries and passages by encoding entire input sequences jointly. Despite these advancements, processing long documents with LLMs remains challenging, primarily due to the quadratic complexity of self-attention mechanisms, leading to high computational and memory demands. Furthermore, there remains room for improvement in the effectiveness of directly using LLMs.
In practical applications, such as legal document retrieval, enterprise search, and academic research, efficiently handling long documents is essential, as relevant information can be spread across lengthy text. Techniques like block selection, which divides long documents into smaller segments and focuses on the most pertinent portions, have emerged as practical solutions to reduce computational demands. The KeyB approach (Li et al., 2023a) has demonstrated that selectively processing key blocks can enhance both retrieval efficiency and accuracy. However, as LLM architectures continue to evolve, revisiting block selection strategies is crucial to fully exploit the potential of these strategies in long document IR.
Moreover, the internal mechanisms of LLMs during the ranking process are still not fully understood, particularly how these models weigh and align relevant content between queries and documents. Understanding the role of attention mechanisms in relevance judgement remains a key challenge, and it is essential to investigate how specific attention heads contribute to aligning query terms with document content in the context of IR.
In this paper, we propose KeyB2, an enhanced block selection strategy designed to integrate seamlessly with LLMs like Llama, combining efficient block selection with the advanced semantic matching capabilities of modern models. Additionally, compared to KeyB models, we incorporate a bi-encoder-based block selection strategy, enhancing the balance between efficiency and semantic comprehension. The KeyB2 approach aims to address the computational challenges associated with long document processing while leveraging LLMs’ contextual understanding to improve ranking performance. By selectively identifying and processing key blocks, KeyB2 reduces computational overhead and enhances retrieval effectiveness, making it well-suited for large-scale IR tasks.
To explore the potential of the KeyB2 approach, we investigate the following research questions:
RQ1:
What occurs within LLMs, such as RankLLaMA, during relevance judgements in IR tasks, and how do attention mechanisms influence the judgement of query and document content?
RQ2:
How can the KeyB approach be upgraded using LLMs, and what are the best strategies for designing effective block selection methods?
RQ3:
Can the integration of block selection strategies with LLMs improve both the efficiency and effectiveness of long document ranking compared to models that process entire documents? Is block selection still crucial in the era of advanced LLMs for long document retrieval, and does it provide an advantage over directly processing entire documents?
We validate our approach with extensive experiments across several long-document datasets, including TREC 2019 DL document ranking task, Robust04, and MLDR-zh. Our findings show that KeyB2 significantly outperforms baseline models, including traditional and LLM-based approaches, in terms of retrieval accuracy. Besides, KeyB2 models have higher computational efficiency compared with LLM based IR approach RankLLaMA. The results highlight that combining LLMs with intelligent block selection can optimize long document ranking while maintaining or improving retrieval performance.
The structure of this paper is as follows: Section 2 reviews related work in IR, LLMs, and block selection methods. Section 3 provides a detailed analysis of attention mechanisms within LLMs and their relevance to document retrieval. Section 4 introduces the KeyB2 approach, while Section 5 outlines the experimental setup, and presents the results of effectiveness and efficiency with analysis. Lastly Section 6 concludes the paper.
2. Related Work
2.1. From Neural-IR Models to Pre-trained Language Models for Information Retrieval
Pre-trained language models (PLMs) have reshaped the field of IR. Initially, models such as the Deep Structured Semantic Model (DSSM) [60] laid the groundwork by employing fully connected networks for representations, mapping queries and documents into a shared semantic space through non-linear projections. Other studies in this area have aimed to leverage distributed representations through variations of DSSM or employed alternative representation functions (Onal et al., 2018; Hu et al., 2014; Shen et al., 2014; Palangi et al., 2016).
Early neural IR models like DRMM (Guo et al., 2016) outperformed traditional methods by utilizing interaction functions and term-level relevance networks. KNRM Xiong et al. (2017) employs a translation matrix using word embeddings to represent word-level similarities. KNRM’s unique kernel-pooling technique extracts multi-level soft match features, and its learning-to-rank layer integrates these features into a final ranking score. This end-to-end trained model tunes word embeddings to generate desired soft matches, demonstrating the effectiveness of neural networks in capturing nuanced relationships between query terms and document content. Inspired by the human judgment process for document relevance, Pang et al. (2017) introduce DeepRank, which splits documents into term-centric contexts for each query term. DeepRank constructs a tensor incorporating the word representations of the query and query-centric context along with their interactions. This tensor is then processed through a measurement network, often based on CNN (Krizhevsky et al., 2012) or 2D-GRU (Wan et al., 2016), to produce a representation of local relevance. The global relevance score is ultimately calculated using an aggregation network, showcasing the ability of PLMs to distill complex textual information into actionable scores. Hui et al. (2017) proposed PACRR, a model influenced by neural models used in image recognition. PACRR uses a similarity matrix between a query and a document as input, with multiple CNN kernels to capture query-document interactions. Building upon PACRR, Hui et al. (2018) developed CO-PACRR, a lightweight contextualization model that averages word vectors within a sliding window, appending the similarities to those of the PACRR model. These developments highlight the versatility of PLMs in adapting to diverse IR requirements.
Transformer-based models, such as BERT (Devlin et al., 2019), have been particularly effective. They leverage attention mechanisms, pre-trained on large scale textual data, enabling a deeper understanding of the semantics embedded within queries and documents. The success of BERT in various NLP tasks, including IR, can be attributed to its capacity to obtain contextualized understanding that account for the surrounding text. It can be classified into three categories of BERT-based IR methods: interaction-based approach, representation-based approach and late-interaction based approach. The first kind is also called cross-encoder, where the query and passage are concatenated and inputted to the BERT model. Nogueira and Cho (2019) proposed using BERT as a cross-encoder for passage re-ranking by fine-tuning the model, achieving significantly higher results than earlier neural-IR models.
The second kind is also called dense retrieval models or bi-encoder. Karpukhin et al. (2020) propose DPR for open-domain question answering, showing simply fine-tuning the question and passage encoders on existing question-passage pairs is sufficient to greatly outperform BM25. They also investigate various negative sampling strategies. Similarly, Reimers and Gurevych (2019) propose Sentence-BERT, which use siamese and triplet network structures to derive semantically meaningful sentence embeddings. Later works focus more on the hard negative mining or domain adaptation for the bi-encoders, such as ANCE (Xiong et al., 2020), SimANS (Zhou et al., 2022) and DoDress (Li and Gaussier, 2024).
ColBERT (Khattab and Zaharia, 2020), a representative approach of the third category, introduces a late interaction architecture that independently encodes each token in the query and the document using BERT, and then employ a cheap yet powerful interaction step that models the fine-grained similarity. Later, ColBERTv2 (Santhanam et al., 2022) is proposed that employs a simple combination of distillation from a cross-encoder and hard-negative mining to boost quality. Then a residual compression mechanism is used to reduce the storage space of the representations.
2.2. Long Document Re-ranking
Dealing with long documents presents a significant challenge for Transformer-based models due to their quadratic complexity. Previously, this complexity arises from the self-attention mechanism, which becomes computationally prohibitive as input length increases, as well as the design of positional embeddings. To overcome these limitations, researchers have explored various strategies to enable Transformer-based models to process long documents effectively. For instance, Dai et al. (2019) proposed a model that incorporates left-to-right recurrence across transformer windows, featuring a segment-level recurrence mechanism combined with a new positional encoding method. Beltagy et al. (2020) propose the Longformer with an attention mechanism that combines local and global information, making it easy to process documents of thousands of tokens or longer. Child et al. (2019) proposed several sparse factorizations of the attention matrix that reduce its computational complexity from quadratic $O(n^{2})$ to $O(n\sqrt{n})$. Zaheer et al. (2021) introduced BigBird, which integrates local and global attention mechanisms along with random sparse attention to handle long sequences. Kitaev et al. (2020) addressed the challenge of efficient self-attention by computing it only between similar tokens, identified through locality-sensitive hashing.
Similarly, for IR dealing with long documents, one line of research uses modified Transformer architecture. Hofstätter et al. (2020) introduced TKL, which employs local self-attention by using a sliding window over the document and limiting the attention to that window, addressing the challenge of processing long documents. Jiang et al. (2020) proposed the Query-Directed Sparse Transformer (QDS-Transformer), which introduces sparsity into the self-attention mechanism by combining local windowed attention with query-directed global attention. Experimental results demonstrate that the QDS-Transformer consistently outperforms previous methods, showing robust advantages across various tasks. However, these modified attention mechanisms still suffer from lower effectiveness and often require custom techniques or CUDA kernels for efficient sparse multiplication on GPUs (Gray et al., 2017; Zaheer et al., 2021; Jiang et al., 2020).
Another line of approach involves breaking long documents into segments, known as passages, and independently processing these segments. Dai and Callan (2019) propose segmenting documents into short, overlapping passages and using BERT to calculate either the first passage’s relevance score, the max score of the passages, or the sum of all passage scores. However, it either lose information or is a little coarse-grained. Later, Li et al. (2023b) propose PARADE model, which utilizes Transformer encoder layers, to combine query-passage relevance signals obtained with BERT, for query-document relevance score. The model shows significant improvements compared with baselines. However, it suffers from high computational costs with high GPU memory requirement, as it need to combine the block level signals meanwhile to the Transformer layers, and each query-block pair is processed by BERT model.
One step further was achieved by approaches aiming to select key blocks. Hofstätter et al. (2021) propose the IDCM approach. Firstly a lightweight Efficient Student Model (ESM) filters out non-relevant passages, and then a more expensive Effective Teacher Model (ETM), like BERT, evaluates the selected passages separately. The ESM is trained using knowledge distillation. Final query-document relevance scores are calculated using a weighted sum of passage scores. The ICLI model (Li and Gaussier, 2022) extends the late-interaction approach to long document retrieval by obtaining a passage representation for each passage in a long document with multi-task learning. Then the representation is used fast intra-passage ranking. Late interaction is then applied to the top-ranking passages to derive document relevance scores. Experimental results demonstrate the effectiveness. Li et al. (2023a) propose KeyB approach, also addresses long document re-ranking by first applying local pre-ranking using traditional IR scoring functions like BM25, or learned scoring methods. The top relevant blocks are then aggregated and processed by a BERT-based model. Unlike IDCM, KeyB processes concatenated blocks as a unified input and shows better performance on datasets like TREC 2019 DL. KeyB allows for a flexible integration with existing models, enhancing scalability. In this paper, we focus on the interaction-based KeyB approach and want to incorporate it with LLMs to see whether in LLMs era, the selecting key block approach is also helpful and can improve the results.
2.3. Reranking with Large Language Models
The utilization of LLMs for reranking has emerged as a promising direction in IR. Unlike earlier models, LLMs, with their vast parameter counts, can potentially grasp the nuances of documents and complex queries. Sun et al. (2023) first investigate prompting ChatGPT with two existing strategies (Sachan et al., 2022; Liang et al., [n. d.]) but observe limited performance. Then they propose an alternative instructional permutation generation approach, instructing the LLMs to directly output the permutations of a group of passages, with a sliding window strategy to address context length limitations. Results demonstrate that GPT-4, equipped with zero-shot instructional permutation generation, surpasses supervised baselines across nearly all datasets.
The RepLLaMA and RankLLaMA models (Ma et al., 2024), exemplify the potential of LLMs in IR. They leverage LLMs directly to serve as components in the widely used multi-stage text ranking pipeline. They fine-tune the open-source Llama-2 model as a dense retriever (RepLLaMA) and a pointwise reranker (RankLLaMA). Through a straightforward training regime, these models exhibit strong zero-shot effectiveness and show that finetuned LLM retrieval models outperform smaller models. Their results show that compared to directly prompting LLMs to generate permutations (Sun et al., 2023), fine-tuning LLMs as rerankers can be more effective. However, in terms of document ranking task, e.g., TREC 2019 DL document ranking, the results do not surpass the BERT-based KeyB approach (Li et al., 2023a). Zhang et al. (2024) propose a two-stage progressive paradigm to adapt LLMs to text ranking. The first stage continuously pretrains LLMs with relevant text pairs on large weakly supervised corpus. The second stage enhances text ranking performance using supervised finetuning and optimized loss functions. Liu et al. (2024) propose PERank, leveraging the single passage embedding as a good context compression for efficient listwise passage reranking. They directly input passage embeddings into LLMs and employ listwise learning to rank loss for training. Experimental results demonstrate that it achieves comparable ranking performance to uncompressed methods while improving inference efficiency. However, this approach requires obtaining passage embeddings and dynamically adjusting the decoding space, which increases architectural complexity and memory usage.
In conclusion, most of these approaches focus on passage retrieval tasks rather than on long-document ranking. RankLLaMA experiment on TREC 2019 DL document ranking task, but the results are lower than KeyB with BERT model. In this paper, we naturally follow the standard RankLLaMA approach and want to see if selecting key block approach is also important in the LLM era and can enhance the results.
3. Attention Mechanisms in LLMs for Information Retrieval
With the advent of PLMs like BERT, IR systems have seen substantial improvements in document ranking accuracy. Among these, re-ranking models, often referred to as cross-encoders, harness the power of fine-tuned PLMs or decoder-only LLMs for downstream IR tasks. Despite the impressive performance of decoder-only LLMs like RankLLaMA, the inner workings of these LLMs, specifically how they assess and rank the relevance of passages, remain underexplored.
To shed light on this, we propose analyzing the attention heatmaps of the rankLLaMA model, which has been fine-tuned on the MSMARCO passage ranking dataset111available at: https://huggingface.co/castorini/rankLlama-v1-7b-lora-passage. A similar investigation by (Clark et al., 2019), even though on a different architecture with less layers and attention heads reveals that a significant portion of BERT’s attention is focused on the delimiter token [SEP], which often attends broadly over entire sentences. In their work, they also observe that certain attention heads align well with linguistic features such as syntax and coreference. However, recent LLMs are uni-directional and decoder-only, and may display a different behavior regarding how tokens attend to each other. We used the BertViz tool (Vig, 2019) to explore these attention patterns in the decoder-only LLM RankLLaMA model when processing various query-passage pairs.
(a) Attention to delimiters
(b) More attention to current tokens
(c) More attention to current tokens
(d) Broad attention
(e) ¡s¿ token exhibits broad attention
(f) ¡/s¿ token exhibits broad attention
(g) Broad attention
(h) ¡s¿ token exhibits broad attention
(i) ¡/s¿ token exhibits broad attention
Figure 1. Attention heads that attend to delimiters, current tokens or attend broadly.
(a) The overall attention map
(b) The attention map for a specific token
(c) The attention map for a specific token
(d) The overall attention map
(e) The attention map for a specific token
(f) The attention map for a specific token
Figure 2. Example attention map of two heads, that potentially are capturing the relevance information.
(a) The overall attention map
(b) The attention map for a specific token
(c) The attention map for a specific token
(d) The overall attention map
(e) The attention map for a specific token
(f) The attention map for a specific token
Figure 3. Attention map of two heads that deal with relevant and irrelevant information. To help visualization, top 3 images are sharpened.
(a) The overall attention map
(b) The attention map for a specific token
(c) The attention map for a specific token
(d) The attention map for a specific token
Figure 4. An example attention map of a head, that captures more relevant information.
3.1. Analysis
We begin with a simple and arbitrary text pair: the query text is ”where is Paris”, and the document text is ”Paris is a city in France”. So the input text to the LLM is ”¡s¿query: where is Paris¡s¿document: Paris is a city in France¡/s¿”. Our findings are summarized in the following points:
  • Attention to delimiter, current and broad tokens: Similar to (Clark et al., 2019) in BERT, we observe that a substantial portion of attention is directed towards the delimiter token ¡s¿, e.g., Fig. 1(a). Besides, we also observe a small amount of attention is directed towards the current token (Fig. 1(b), Fig. 1(c)). The authors in (Clark et al., 2019) speculate that attention over the ¡s¿ delimiter tokens might be used as a sort of “no-op”, and we conjecture that the attention over the beginning ¡s¿ and current tokens is also ”no-op” for IR scenario. We also observe that some attention heads attend broadly over all tokens (Fig. 1(d), Fig. 1(g)), which is consistent with (Clark et al., 2019).
    The delimiter tokens, as ¡s¿ and ¡/s¿, are important in PLMs and LLMs as, respectively the CLS token, separation token or the last delimiter token, are used to separate the input or compute the retrieval status value (RSV) of a document to a given query. It is thus important to examine what happens with these delimiter tokens, and in particular to the last delimiter for Llama-based models. We can see in Fig. 1(e) that the second token ¡s¿ attends broadly over the query tokens, and on Fig. 1(f) that the token ¡/s¿ attends broadly over both the query and the document tokens. This suggests that these delimiters capture the overall representations of, respectively, the query and of the query-document pair. In Fig. 1(h) and Fig. 1(i), they also perform similarly but attend more to salient tokens, in the query for the second ¡s¿ token and in the document for the token ¡/s¿. As the last delimiter ¡/s¿, at the basis of the RSV calculation in RankLLaMA, also exhibits broad attention, it can eventually be viewed as a coarse-grained representation of the query-document pair dedicated to IR.
  • Attention between relevant tokens: Interestingly, when examining other attention heads, we find that several heads seem to capture relations between relevant tokens in the query and document. As shown in Fig. 2, for Attention Head 25 in Layer 1, the token “Paris” in the document attends primarily to “Paris” in the query, while the token “France” in the document attends to “Paris” in both the query and document. For Attention Head 25 in Layer 8, tokens in the document tend to attend to relevant tokens in the query. These illustrate that specific attention heads are adept at identifying relevant matches between query terms and document content, which is essential for accurate relevance judgement.
  • Attention with irrelevant information: To test what happens between irrelevant tokens, we introduce irrelevant information, such as “An apple is a round fruit”, into the document. As shown in Fig. 3, the attention head 25 in layer 8 suppresses the irrelevant content, maintaining strong attention on the relevant tokens “Paris” and “France.” This head is trying to discard the noise (e.g., apple) and focuses on query-relevant portions. However, head 25 in layer 1, which displays broad attentions over all tokens, shows that the attention connections can be more contrasted. Fig. 3(b) shows that the token “France” attends more to Paris than to other tokens, and that the last token ¡/s¿ attends to both relevant tokens, “Paris” and “city”, as well as irrelevant tokens like “apple”, “is”, “a”, although weakly. These irrelevant tokens will inevitably result in noisy information in the representation of the last token “¡/s¿”, which is the basic building block for computing the relevance score for RankLLaMA.
  • Dynamic adjustment with additional relevant information: We further analyze how the model handles additional relevant information by expanding the document to include another sentence about Paris, as shown in Fig. 4 (to be compared with Fig. 3). The attention map shows that the attention head tries to adjust its focus to cover both instances of relevant content. This is similar to the way humans construct relevance judgments for long documents (Wu et al., 2007), by assessing the relevance of blocks within documents.
3.2. Discussion
These insights reveals part of the internal processes behind RankLLaMA. The attention mechanisms in Llama models, as explored in this section, show similarities to the way humans assess the relevance of documents, by selectively focusing on relevant information while filtering out irrelevant content. For example, given a query like ”Where is Paris?” in a passage such as ”Paris is a city in France. An apple is a round fruit,” humans naturally prioritize the tokens ”Paris” and ”France,” while ignoring irrelevant parts like ”apple.” Similarly, Llama models are trying to concentrate their attention on query-relevant terms, as seen in Fig. 2, Fig. 3 and Fig. 4. However, certain attention heads attend broadly, as shown in Fig. 3(a) and Fig. 3(c), and the last token ¡/s¿ weakly attends to irrelevant tokens, resulting in noisy representations and potentially erroneous relevance scores.
Although recent LLMs can handle input sizes from thousands to hundreds of thousands of tokens, their computational complexity still depends on the input token count. Similar to the Transformer or BERT architectures, recent LLMs like Llama2 or Llama3 also use self-attention mechanism. The difference is recent LLMs use the decoder-only architecture with unidirectional attention. Llama models use Grouped Query Attention (GQA) (Ainslie et al., 2023): instead of each attention head having its own key (K) and value (V) matrices, multiple heads are grouped together, and each group shares the same key and value vectors. This mechanism is designed to reduce memory and computational overhead, particularly in large models like Llama and Mistral (Jiang et al., 2023), by allowing several query heads to reuse the same key-value pair within a group. This helps speed up inference while maintaining the benefits of multi-head attention. Recent LLMs also use KV Cache (Pope et al., 2023) to speed up inference, but the KV Cache still faces memory bound issue. Besides, the cache does not apply to RankLLaMA. Indeed, the KV Cache technique caches the Key (K) and Value (V) matrices from previous tokens, enabling the model to reuse these stored results when generating new tokens. This avoids redundant computations and significantly improves efficiency. However, the RankLLaMA model only use the ¡/s¿ token to infer the relevance once and does not need to generate more tokens. Based on the above analyses, the complexity of the RankLLaMA model remains quadratic relative to the input token count.
The block selection strategy for IR (Li et al., 2023a) mimics human filtering by choosing the most relevant passages or blocks from long documents, reducing computational load and improving retrieval performance by focusing on critical content and removing noisy contents. Pre-ranking blocks enhances both the efficiency and accuracy of relevance judgements with BERT. Similarly, for LLMs used in IR, such as RankLLaMA, a pre-ranking strategy for passages or blocks may help models avoid irrelevant tokens and the noisy representations they may lead to, and improve both efficiency and effectiveness in long document retrieval tasks. This is the question we are addressing now.
4. Method: KeyB2
Figure 5. The architecture of proposed KeyB2 approach.
Based on the above conlusion, we show in this paper how to apply to LLMs the KeyB block selection approach proposed in (Li et al., 2023a) for long document IR. This answers our second research question regarding the design of block selection methods for LLM based IR rankers. The systems we propose, illustrated in Figure 5, are based on block selection mechanisms using fast models as standard IR models or PLMs, followed by LLM rankers based on Llama2 and Llama3, using a similar generation strategy as RankLLaMA.
This section introduces the long document segmentation approach, followed by three distinct block pre-ranking methods, including a novel one, and concludes with an aggregation of the top-scoring blocks.
4.1. Passage Segmentation Approach
When reading a document for assessing its relevance with respect to a given query, humans tend to segment the document into pieces prior to scan each piece for assessing its relevance (Ding et al., 2020). Inspired by this, passage or block segmentation methods have become popular in IR. We use in this paper the CogLTX (Ding et al., 2020) block decomposition method, which have been applied with success on several NLP tasks and which sets different costs for different punctuation marks with the aim to segment in priority on strong punctuation marks such as ”.” and ”!”, making it close to the sentence segmentation procedure used by humans. CogLTX segments each long text $D$ into a sequence of blocks $[b_{1}\dots b_{n}]$ using dynamic programming. This process ensures that the length of each block does not exceed a predefined maximum value B; in our implementation, B=63, which is the same as the setting in (Li et al., 2023a).
Furthermore, to be consistent with the final LLM-based IR ranker, we use different tokenizers depending on the IR ranker used: the Llama2 tokenizer is used when the IR ranker is Llama2 whereas the Llama3 tokenizer is used when the IR ranker is Llama3. Last but not least, for Chinese documents, Since in Chinese, marks imported from Europe are fullwidth instead of halfwidth like their original European counterparts [^2] https://en.wikipedia.org/wiki/Chinese_punctuation. So in Chinese, the punctuations are different with English in terms of: comma, period, exclamation mark and question mark, and we incorporate various Chinese punctuation marks into CogLTX. This ensures more effective separation of Chinese sentences.
4.2. Local ranking approaches for KeyB2
After segmenting the long document into blocks, the keyB2 model efficiently calculate the local relevance. This step should be significantly faster than using LLMs. The IR community has evolved in the past decades, from bag of words approaches to using neural networks including Transformer based networks. Different IR approaches can be plugged in the KeyB2 model. In this paper, we adopt three different approaches: they are BM25, cross-encoder and bi-encoder models. These three approaches are widely used in academia and industries.
4.2.1. KeyB2-BM25
BM25 is based on term matching mechanism. It improves TF-IDF approach which is also based on term matching. However, they adds some statistical information by calculating the inverse document frequency (IDF). Besides, compared with TF-IDF, BM25 introduces more parameters to modulate the impact of term frequency and takes into account the effects of document length and query length. Therefore, its performance is typically superior to that of TF-IDF in practical applications.
Given a query and a block, the Retrieval Status Value (RSV) is calculated as below:
$RSV(q,b)_{BM25}=\sum_{w\in q\cap b}IDF(w)\cdot\frac{tf_{w}^{b}}{k_{1}\cdot(1-b+b\cdot\frac{l_{b}}{l_{avg}})+tf_{w}^{b}},$
in which $tf_{w}^{b}$ corresponds to the number of occurrences of word $w$ in block $b$ and $IDF(w)$ to the inverse document frequency of word $w$. In this paper, we use the scikit-learn (Pedregosa et al., 2011) package to calculate the IDF dictionary with IDF smoothing. $IDF(w)$ is defined by:
$IDF(w)=\log\frac{N+1}{df_{w}+1}+1,$
where $N$ and $df_{w}$ are added one with default scikit-learn smooth IDF setting. The added 1 after the fraction, makes sure terms with zero IDF don’t get suppressed entirely. Besides, $l_{b}$ is the length of block $b$, $l_{avg}$ the average length of the blocks in $d$, and $k_{1}$ and $b$ two hyperparameters.
For Chinese documents, word $w$ is recognized using the Jieba [^3] https://github.com/fxsjy/jieba Chinese text segmentation tool to obtain meaningful terms.
4.2.2. KeyB2-cross
For each block $b_{i}$, to calculate its relevance with the query, we also can use a cross-encoder model, which should be efficient since the length of each block is limited. since the block number is relatively large, this model should be relatively fast. We use pre-trained language models (PLMs) like BERT (Devlin et al., 2019) as the cross-encoder, with a typical number of parameters ranging from dozens of millions to hundreds of millions, which is one to two orders of magnitude smaller than the one of LLMs. Denoting by $b_{i}$ the $i^{th}$ block and $b_{i}^{cls}$ the corresponding query-block representation, one has:
$b_{i}^{cls}=\text{PLM}([CLS],q\_tokens,[SEP],b_{i}\_tokens),$
where PLM is a pre-trained language model like BERT, $q\_tokens$ are the tokens of the query and $b_{i}\_tokens$ are the tokens of the block. A linear layer is then used to generate the query-block relevance score:
$RSV(q,b_{i})=Wb_{i}^{cls},$
where $W$ is the learnable weights of a feedforward neural network.
4.2.3. KeyB2-bi
The previous subsection uses a cross-encoder to identify the query-block relevance. Although cross-encoders have shown their effectiveness, they have the drawback of high computational latency since the concatenated query-block tokens should be calculated online. On another side, bi-encoders are another kind of neural-IR approaches which are very efficient because the block representations can be calculated offline (also called dense retrieval). Besides, compared with BM25, they have the ability to capture the semantic relevance information. Normally, the query encoder and passage/block encoder can be shared, and in this paper, we follow this paradigm: a same encoder is used for queries and blocks.
Given a query or a block, the bi-encoder model encodes the tokens of the query or the tokens of the block into a high dimensional representation:
$q^{cls}=\text{PLM}([CLS],q\_tokens),$
$b_{i}^{cls}=\text{PLM}([CLS],b_{i}\_tokens).$
Then the retrieval status value is calculated by either:
$RSV(q,b_{i})=q^{cls}\cdot b_{i}^{cls},$
or
$RSV(q,b_{i})=\cos(q^{cls},b_{i}^{cls}).$
4.3. Aggregating the Local Relevance
To aggregate the different blocks deemed relevant to the query, the KeyB2 approach selects the most relevant blocks given a query and aggregate the tokens of these blocks with the query in order to generate the overall document relevance score using LLMs. More preceisely, given the scores of blocks: $S=[s_{1}\dots s_{n}]$, and the max combined block token limitation $max\_b\_t$, the selected top $k$ blocks from the list $S$ are $B_{r}=[b_{r_{1}},b_{r_{2}},\dots,b_{r_{k}}]$, which satisfy:
(1) $\sum_{i=1}^{k}L(b_{r_{i}})\geq max\_b\_t,$
and
(2) $\sum_{i=1}^{k-1}L(b_{r_{i}})\textless max\_b\_t.$
The tokens for the newly concatenated document are:
(3) $Sort_{index}(B_{r})=[b^{1}\dots,b^{i},\dots,trunc(b^{k})],$
where $b^{i}$ is the $i\_th$ block in the top $k$ scoring blocks $B_{r}$, after having re-ranked the list according to their original indexes in the original document, and $trunc(b^{k})$ means the last block is truncated according to $max\_b\_t$, similar to (Li et al., 2023a). This step ensures the top $k$ blocks are re-ordered according to their original appeared locations in the document, to allow a coherent meaning and structure.
Finally, the step goes to calculating the RSV between the query and $Sort_{index}(B_{r})$, in this paper, the KeyB2 model follows RankLLaMA approach to calculate the relevance score. The formated input tokens are:
[…]
and the RSV is:
(4) $RSV(Q,D)=\text{Linear}(\text{Decoder}(\text{input})[-1])$
5. Experiment
In this section, we evaluate the proposed approach on various datasets and compare it with baseline models. This section mainly aims at addressing RQ2 and RQ3 as introduced in Section 1.
5.1. Datasets
Since this paper focuses on document ranking, the datasets should be relatively long. In this paper, we use 3 datasets:
  • The first dataset is TREC 2019 DL document ranking task (Craswell et al., 2020). The Deep Learning Track, introduced for TREC 2019, aims to study large data ad hoc ranking. It is the first track large human-labeled training datasets. It includes two tasks: document ranking and passage ranking. In this paper, we focus on the document ranking task.
  • The second dataset is Robust04444https://trec.nist.gov/data/robust/04.guidelines.html, whose sources include Financial Times, Federal Register 94, FBIS, disk 5 and LA Times.
  • The last dataset MLDR Multilingual Long-Document Retrieval dataset (Chen et al., 2024), and we use the MLDR-zh (Chinese) subset. The source of MLDR-zh are Wikipedia and Wudao (Yuan et al., 2021).
5.2. Baselines
To evaluate the performance of KeyB2 models, we compare them against several baseline approaches across multiple datasets:
  • On TREC 2019 DL: Following (Li et al., 2023a), the baselines used for this dataset include traditional methods like BM25, neural IR models such as CO-PACRR (Hui et al., 2018), TK (Hofstätter et al., 2020), and TKL (Hofstätter et al., 2020), as well as transformer-based models like RoBERTa (FirstP and MaxP) (Liu et al., 2019) and PARADE (Li et al., 2020). Additionally, we compare with sparse attention models, including Sparse-Transformer (Child et al., 2019), Longformer-QA (Beltagy et al., 2020), Transformer-XH (Zhao et al., 2020), and QDS-Transformer (Jiang et al., 2020). Previous key block selection models like IDCM (Hofstätter et al., 2021) and KeyB (Li et al., 2023a) are also included. The LLM-based RankLLaMA (Ma et al., 2024) model serves as another important baseline.
  • On Robust04: The baselines for this dataset consist of BM25, as well as BERT-based KeyB models (using BM25 and BinB configurations). We also compare with the RankLLaMA model.
  • On MLDR-zh: the baselines include traditional IR methods such as BM25, along with KeyB models using BM25 and BinB configurations. For this dataset, the KeyB models utilize a Chinese version of BERT and a BM25 implementation adapted for Chinese text. The RankLLaMA model is also used as a baseline for multilingual retrieval.
5.3. Experimental Design and Implementation Details
Experimental Design
To evaluate KeyB2, we first fine-tuned the models on the TREC 2019 DL document ranking dataset, comparing them with various baselines. For a fair comparison with RankLLaMA, which uses the Llama2555https://huggingface.co/meta-llama/Llama-2-7b-hf model with 7B size, we initially experimented with KeyB2 using the same Llama2, referred to as KeyB2(Llama2). To explore potential improvements, we also tested KeyB2 with Llama3666https://huggingface.co/meta-llama/Meta-Llama-3-8B model with 8B size, named KeyB2(Llama3). The KeyB2 approach includes three key block selection methods: BM25, cross-encoder, and bi-encoder. We fine-tuned KeyB2(Llama2) and KeyB2(Llama3) for each method, resulting in configurations such as KeyB2(Llama2)BM25, KeyB(Llama2)cross and KeyB(Llama2)bi.
Next, we evaluated the zero-shot performance by applying each KeyB2 model, trained on the TREC 2019 DL dataset, to the Robust04 dataset without additional fine-tuning. Finally, we conducted other fine-tuning experiments, where both KeyB2 and baseline models were trained from scratch on the downstream datasets Robust04 and MLDR-zh. The tasks of all models excluding BM25 for the datasets are to rerank documents in the test sets, while BM25 shows the retrieval performances (Yang et al., 2018).
Implementation Details
We implemented KeyB2 and RankLLaMA using PyTorch 2.4.0 and CUDA 11.8, with training conducted on an NVIDIA A100 GPU. The main configurations and training details are as follows:
  1. (1)
    Input Size Settings:
    • The maximum query token number is 32. For KeyB2, the document token number is limited to 480 after block selection, resulting in a total input size of 512 tokens, matching the KeyB (Li et al., 2023a) limit.
    • For RankLLaMA, which processes documents directly, the maximum token number is set to 1200 to fit within A100 GPU.
  2. (2)
    Training Configurations:
    • For KeyB2 models, the batch size is set to 2. To efficiently fine-tune large language models (LLMs), we employ the LoRA technique (Hu et al., 2021), which freezes LLM weights and introduces trainable matrices to reduce the number of trainable parameters. LoRA parameters were set to $r=32$ and $\alpha=64$.
    • The learning rate is $5\times 10^{-5}$, with a gradient accumulation step of 8. For RankLLaMA, the batch size is 1 and other settings are the same with KeyB2. AdamW optimizer is used with WarmupDecayLR scheduler using DeepSpeed (Rasley et al., 2020), with FP16 mixed precision.
    • For KeyB models, following the code from (Li et al., 2023a) with the learning rate $5\times 10^{-5}$.
  3. (3)
    Model Variants for Different Languages:
    • KeyB2 uses small pre-trained models for cross-encoders or bi-encoders depending on the dataset language, as detailed in Table 1.
  4. (4)
    BM25 Implementation:
    • For English datasets, BM25 is implemented using the Anserini tool (Yang et al., 2018) with default settings ($k_{1}=0.9$, $b=0.4$). For KeyB2-BM25, we computed the IDF dictionary using scikit-learn on the TREC 2019 DL dataset, with parameters matching Anserini’s defaults.
    • For the Chinese MLDR-zh dataset, the BM25 baseline is also implemented with Anserini, using default settings. For KeyB2-BM25, as mentioned in Section 4.2.1, the sentences are separated into Chinese words. Other settings are the same to English datasets.
  5. (5)
    Training Data:
    • TREC 2019 DL: 200K triplets of a query, a positive document, and a negative document.
    • Robust04: 5-fold cross-validation with 40K triplets per fold, and fold separation follows (Li et al., 2023a).
    • MLDR-zh: 10,000 training queries, each with a positive and negative document.
  6. (6)
    Fine-Tuning Approach:
    • Fine-tuning uses pairwise hinge loss with a learning rate of $5\times 10^{-5}$.
    • Models are initialized from official Llama checkpoints.
  7. (7)
    Evaluation Metrics:
    • On TREC 2019 DL, following previous works (Li et al., 2020, 2023a), NDCG@10 and MAP are used.
    • The evaluation metrics include precision at 1, 5, 10, and 20 (P@k), as well as NDCG@k, NDCG and MAP (Li et al., 2023a).
    • On MLDR-zh, in the qrel file of MLDR Chinese subset, each query corresponds to one positive label, and in the test file, each query consists 8 documents for ranking. So for the P@k metric, we report the P@1, indicating the percentage of the relevant document is ranked in the first position by the approaches. Besides, MAP is reported. For NDCG@k, we report values at NDCG@1, NDCG@5, and NDCG@8.
Table 1. The cross-encoders and bi-encoders that are used in KeyB2 models for datasets with different languages (huggingface Transformer versions).
TREC 2019 DL & Robust04 MLDR-zh
cross-encoder cross-encoder/ms-marco-MiniLM-L-6-v2 BAAI/bge-reranker-base
bi-encoder sentence-transformers/all-MiniLM-L6-v2 DMetaSoul/Dmeta-embedding-zh-small
5.4. Fine-tuning on TREC DL19 Document Ranking Task
Table 2. Experiment on TREC 2019 DL and comparison with sparse attention models and previous select blocks models. Baseline results before the LLM and KeyB2 parts are from (Li et al., 2023a). Best results are in bold.
TREC Deep Learning Track Document Ranking
Model NDCG@10 MAP
Baseline models
BM25 0.488 0.234
CO-PACRR (Hui et al., 2018) 0.550 0.231
TK (Hofstätter et al., 2020) 0.594 0.252
TKL (Hofstätter et al., 2020) 0.644 0.277
RoBERTa (FirstP) (Liu et al., 2019; Dai and Callan, 2019) 0.588 0.233
RoBERTa (MaxP) (Liu et al., 2019; Dai and Callan, 2019) 0.630 0.246
PARADE (Li et al., 2020) 0.655 0.280
Sparse attention models
Sparse-Transformer (Child et al., 2019) 0.634 0.257
Longformer-QA (Beltagy et al., 2020) 0.627 0.255
Transformer-XH (Zhao et al., 2020) 0.646 0.256
QDS-Transformer (Jiang et al., 2020) 0.667 0.278
Select blocks models
IDCM (Hofstätter et al., 2021) 0.679 0.273
KeyB(BERT)BM25 (Li et al., 2023a) 0.678 0.277
KeyB(BERT)BinB (Li et al., 2023a) 0.707 0.281
KeyB(PARADE5)BM25 (Li et al., 2023a) 0.672 0.280
KeyB(PARADE5)BinB (Li et al., 2023a) 0.676 0.277
KeyB(PARADE5)BinB2 (Li et al., 2023a) 0.678 0.279
LLM
RankLLaMA-original paper (Ma et al., 2024) 0.677 -
RankLLaMA-trainByUs1200 0.707 0.291
KeyB2 models
KeyB2(Llama2)BM25 0.713 0.293
KeyB2(Llama2)cross 0.730 0.302
KeyB2(Llama2)bi 0.710 0.293
KeyB2(Llama3)BM25 0.718 0.293
KeyB2(Llama3)cross 0.731 0.298
KeyB2(Llama3)bi 0.711 0.296
The TREC DL19 document ranking task is representative for long document IR experiments as it includes a large amount of training data suitable for effective model training. The results demonstrate that the KeyB2 models outperform multiple baseline approaches on the TREC 2019 DL document ranking task, highlighting the effectiveness of the proposed approach. KeyB2 models, particularly those using Llama2 and Llama3 with cross-encoder configurations, consistently achieve higher NDCG@10 and MAP scores than baseline methods, sparse attention models, and prior key block selection techniques.
Performance Highlights:
  • KeyB2 vs. Traditional and Neural IR Baselines: The KeyB2 models achieve largely gains over traditional methods like BM25 and neural IR models such as CO-PACRR and PARADE. For example, KeyB2(Llama3)cross improves NDCG@10 by more than 30% over BM25 and by approximately 12% over PARADE. This indicates that the block selection strategy in KeyB2 effectively enhances the ranking of long documents by focusing on the most relevant content.
  • Comparison with Sparse Attention Models: KeyB2 outperforms sparse attention models such as Longformer-QA and QDS-Transformer, with NDCG@10 improvements of around 10%. This suggests that selecting key blocks is more effective than applying sparse attention mechanisms to the entire document.
  • Advantage over Previous Key Block Selection Methods: Compared to earlier key block selection models like KeyB(BERT), KeyB2(Llama3)cross increases the result in NDCG@10 from the best 0.707 to 0.731, and KeyB2(Llama3)BM25 increases the NDCG@10 from 0.678 to 0.718. The results also largely outperform KeyB(PARADE5). These demonstrate the benefit of integrating LLMs with the block selection approach.
  • KeyB2 vs. RankLLaMA: With the same Llama model, The KeyB2(Llama2)cross configuration improves the score in NDCG@10 from 0.677 (reported in (Ma et al., 2024)) to 0.730, or 0.707 (trained by us using pairwise Hinge loss) to 0.730. This indicating that focusing on selected blocks leads to better ranking accuracy than processing whole documents directly.
Key Observations:
The findings suggest that combining LLMs with a key block selection strategy largely enhances retrieval performance for long document ranking tasks. This answers RQ3. The improvements with cross-encoder configurations indicate that leveraging semantic-rich representations for relevance matching is more effective, partly answers RQ2. The consistent gains across different KeyB2 settings confirm the robustness and adaptability of the proposed approach.
In terms of Llama2 and Llama3, on this dataset, it shows similar results.
5.5. Zero-Shot Performance on Robust04
We load the checkpoint of each model trained in Section 5.4 and evaluate each model on Robust04.
Results: As shown in Table 3, the previous KeyB(BERT)BM25 model performs worse than BM25 on average, whereas the KeyB2 models maintain competitive performance without additional training on Robust04. The KeyB2 models all demonstrate superior zero-shot performance compared to RankLLaMA. Notably, KeyB2(Llama3)cross achieved the highest scores, with NDCG@20 of 0.4887 and P@20 of 0.4187, followed closely by KeyB2(Llama3)BM25, which reached NDCG@20 of 0.4758 and P@20 of 0.4129. In comparison, RankLLaMA recorded NDCG@20 of 0.4472 and P@20 of 0.3873. These results indicate that the KeyB2 models, especially the Llama3-based configurations, generalize effectively to unseen datasets.
Significance Testing: We perform significant tests on the Robust04 dataset, comparing KeyB2 models with RankLLaMA and previous BERT-based KeyB models. The results indicate that KeyB2 models using the cross-encoder block selection strategy significantly outperformed RankLLaMA across all metrics. Over half of the metrics for BM25 or bi-encoder strategies also showed significant improvement over RankLLaMA. However, when compared to KeyB(BERT)BinB, only the KeyB2(Llama3)cross configuration displayed significant improvements across most metrics. When comparing KeyB2 models using Llama3 with using Llama2, only KeyB2(Llama3)cross show significant differences as well.
The zero-shot evaluation results demonstrate that KeyB2 models, particularly those utilizing the Llama3 model with cross-encoder strategies obtains the best results. While the significant tests of other KeyB2 models, suggest further investigation might be needed. This leads to the question of whether fine-tuning on Robust04 might yield better findings.
Table 3. Zeroshot Results on Robust04 dataset. Best results are in bold. A paired t-test ($p-value\leq 0.05$) is used for measuring significance. For KeyB2 models, a significant difference with RankLLaMA is marked with a ’R’, with KeyB(vBERT)BinB with a ’B’. We also report the comparisons between the Llama2 and Llama3 in KeyB2 models. If a KeyB2-Llama3 model is significant different to a KeyB2-Llama2 model, with the same block selection approach, it is marked with †.
Model P@1 P@5 P@10 P@20 MAP NDCG@1 NDCG@5 NDCG@10 NDCG@20 NDCG
Baseline models
BM25 0.5542 0.5004 0.4382 0.3631 0.2334 0.5080 0.4741 0.4485 0.4240 0.4402
BERT based models
KeyB(BERT)BM25 0.5502 0.4787 0.4368 0.3716 0.2226 0.5080 0.4652 0.4461 0.4258 0.4196
KeyB(BERT)BinB 0.6107 0.5269 0.4699 0.3958 0.2444 0.5705 0.5087 0.4845 0.4606 0.4369
LLM
RankLLaMA 0.5663 0.5100 0.4518 0.3873 0.2296 0.5241 0.4904 0.4646 0.4472 0.4269
KeyB2 models
KeyB2(Llama2)BM25 0.5662 0.5357R 0.4880R 0.4074R 0.2404R 0.5341 0.5161R 0.4981R 0.4713R 0.4356R
KeyB2(Llama2)cross 0.6265R 0.5349R 0.4739R 0.4020R 0.2406R 0.5843R 0.5241R 0.4934R 0.4692R 0.4378R
KeyB2(Llama2)bi 0.5984 0.5422R 0.4723R 0.3994R 0.2384R 0.5522 0.5228R 0.4899R 0.4655R 0.4347R
KeyB2(Llama3)BM25 0.6024 0.5390R 0.4888R 0.4129RB 0.2472R† 0.5602 0.5191R 0.5006R 0.4758R 0.4390R
KeyB2(Llama3)cross 0.6305R 0.5622RB† 0.4996RB† 0.4187RB† 0.2520R† 0.5904R 0.5419RB 0.5162RB† 0.4887RB† 0.4462RB†
KeyB2(Llama3)bi 0.5743 0.5213 0.4779R 0.4074R 0.2437R 0.5341 0.5043 0.4895R 0.4693R 0.4381R
5.6. Fine-tuning on Robust04 and MLDR-zh
We now want to compare the results of KeyB2 models after fine-tuning on Robust04, not the zero-shot setting.
Table 4. Fine-tuned models’ results on Robust04 dataset. Best results are in bold. A paired t-test ($p-value\leq 0.05$) is used for measuring significance. For KeyB2 models, a significant difference with RankLLaMA is marked with a ’R’, with KeyB(vBERT)BinB with a ’B’. We also report the comparisons between the Llama2 and Llama3 in KeyB2 models. If a KeyB2-Llama3 model is significant different to a KeyB2-Llama2 model, with the same block selection approach, it is marked with †.
Model P@1 P@5 P@10 P@20 MAP NDCG@1 NDCG@5 NDCG@10 NDCG@20 NDCG
Baseline models
BM25 0.5542 0.5004 0.4382 0.3631 0.2334 0.5080 0.4741 0.4485 0.4240 0.4402
BERT based models
KeyB(BERT)BM25 0.6347 0.5445 0.4887 0.4161 0.2473 0.5924 0.5311 0.5043 0.4805 0.4404
KeyB(BERT)BinB 0.6625 0.5614 0.5007 0.4226 0.2556 0.6243 0.5446 0.5163 0.4902 0.4458
LLM
RankLLaMA 0.5700 0.5412 0.4947 0.4218 0.2508 0.5359 0.5178 0.5005 0.4804 0.4377
KeyB2 models
KeyB2(Llama2)BM25 0.6744R 0.6063RB 0.5272R 0.4447RB 0.2668R 0.6222R 0.5799RB 0.5416R 0.5140RB 0.4521R
KeyB2(Llama2)cross 0.6586R 0.5758R 0.5023 0.4329 0.2554 0.6024R 0.5546R 0.5201 0.4992R 0.4464R
KeyB2(Llama2)bi 0.6504R 0.5959R 0.5348RB 0.4482RB 0.2691R 0.6023R 0.5728R 0.5458RB 0.5168RB 0.4553R
KeyB2(Llama3)BM25 0.7186R 0.6289RB 0.5643RB† 0.4749RB† 0.2822RB† 0.6684R 0.6033RB 0.5755RB† 0.5472RB† 0.4654RB†
KeyB2(Llama3)cross 0.7313R† 0.6514RB† 0.5811RB† 0.4799RB† 0.2923RB† 0.6851R† 0.6302RB† 0.5977RB† 0.5593RB† 0.4718RB†
KeyB2(Llama3)bi 0.6984R 0.6264RB† 0.5665RB† 0.4764RB† 0.2883RB† 0.6502RB 0.6006RB 0.5748RB† 0.5461RB† 0.4654RB†
Results on Robust04: As shown in Table 4, all neural models largely outperform the BM25 results. RankLLaMA obtains lower results than KeyB(BERT)BinB, indicating directly inputting a long document to LLMs might not be a good solution. The proposed KeyB2(Llama3)cross model achieve the best performance on all metrics, with NDCG@20 = 0.5593 and P@20 = 0.4799, surpassing all baseline models, including RankLLaMA with NDCG@20 = 0.4804 and P@20 = 0.4218, the best KeyB model KeyB(BERT)BinB with NDCG@20 = 0.4902 and P@20 = 0.4226. In terms different Llama base models, the KeyB2 using Llama3 obtains overall better results.
Results on MLDR-zh: As shown in Table 5, all models outperform the BM25 baseline by a significant margin. RankLLaMA scores are lower than those of both KeyB models with BERT, suggesting that directly feeding long documents to LLMs may not be the most effective approach for this task. The KeyB2(Llama2)bi configuration obtains the best results across all metrics, with NDCG@8 = 0.9744 and MAP = 0.9661, outperforming all baseline models. The MAP score is higher than RankLLaMA of 115.4%, whose is 0.4485, and higher than the good-performing KeyB(BERT)BinB model of 10.3%, which reaches 0.8758. On MLDR-zh dataset, the bi-encoder strategy looks overall better.
Significance Testing: On fine-tuned Robust04, comparing with RankLLaMA, all models significantly outperform it, with most metrics of KeyB2(Llama2)cross and all metrics of other KeyB2 models. Comparing with the best KeyB model KeyB(BERT)BinB, four metrics of KeyB2(Llama2)BM25 and KeyB2(Llama2)bi significantly improve it, and most metrics of KeyB2(Llama3) models show significant differences. In terms of using Llama2 or Llama3, all metrics of cross-encoder strategy shows significant difference and the other two shows significant improvements with 6 metrics.
On MLDR-zh dataset, all metriecs of proposed KeyB2 models show significant improvements comparing with RankLLaMA and KeyB(BERT)BinB. In terms of comparing models using Llama2 and Llama3, KeyB2 models with BM25 or cross-encoder show significant differences.
Key Observations: On these two datasets, the fine-tuned results show the proposed KeyB2 models can significantly improve the baseline models, including RankLLaMA which directly input long sequences. These results address RQ3, demonstrating the importance of key block selection in the era of LLMs. In terms of different strategies, results show either using cross-encoder or bi-encoder would obtain the best results, on different datasets. All strategies can be successfully incorporated in KeyB2 models, showing the flexibility and adaptability. Comparing with the zero-shot performance, fine-tuning on Robust04 shows larger improvements, and better potential of using newer Llama model.
Table 5. Results on MLDR Chinese dataset. Best results are in bold. A paired t-test ($p-value\leq 0.05$) is used for measuring significance. For KeyB2 models, a significant difference with RankLLaMA is marked with a ’R’, with KeyB(vBERT)BinB with a ’B’. We also report the comparisons between the Llama2 and Llama3 in KeyB2 models. If a KeyB2-Llama3 model is significant different to a KeyB2-Llama2 model, with the same block selection approach, it is marked with †.
Model P@1 MAP NDCG@1 NDCG@5 NDCG@8
Baseline models
BM25 0.2013 0.2591 0.2013 0.2629 0.2768
BERT based models
KeyB(BERT)BM25 0.4225 0.5862 0.4225 0.6249 0.6846
KeyB(BERT)BinB 0.8038 0.8758 0.8038 0.9001 0.9066
LLM
RankLLaMA 0.2975 0.4485 0.2975 0.4439 0.5752
KeyB2 models
KeyB2(Llama2)BM25 0.8638RB 0.9099RB 0.8638RB 0.9231RB 0.9319RB
KeyB2(Llama2)cross 0.9038RB 0.9394RB 0.9038RB 0.9497RB 0.9543RB
KeyB2(Llama2)bi 0.9475RB 0.9661RB 0.9475RB 0.9706RB 0.9744RB
KeyB2(Llama3)BM25 0.9088RB† 0.9420RB† 0.9088RB† 0.9520RB† 0.9563RB†
KeyB2(Llama3)cross 0.9325RB† 0.9584RB† 0.9325RB† 0.9653RB† 0.9688RB†
KeyB2(Llama3)bi 0.9462RB 0.9648RB 0.9462RB 0.9705RB 0.9735RB
5.7. Analysis of the Position of Selected Blocks
Figure 6. The probabilities of top 8 block appearing locations of TREC 2019 DL with different block selecting methods.
Figure 7. The probabilities of top 8 block appearing locations of Robust04 with different block selecting methods.
Figure 8. The probabilities of top 8 block appearing locations of MLDR-zh with different block selecting methods.
We analyze the block selection patterns across different strategies (BM25, cross-encoder, and bi-encoder) and datasets, as shown in Fig. 6, Fig. 7, and Fig. 8. These figures show the probability distributions of the top 8 selected blocks across various document positions, highlighting trends in how different methods prioritize blocks within long documents.
Dataset-Specific Observations
  • TREC DL19 (Fig. 6): All three methods show similar preferences for blocks in the front half. The cross-encoder focuses more on the front half, while the BM25 approach continues to emphasize the later positions, particularly the last three.
  • Robust04 (Fig. 7): While early positions are still preferred, middle and later parts contribute more, particularly with the cross-encoder and bi-encoder methods. These strategies show a more distributed selection, retrieving relevant information from various positions and displaying a more spread-out pattern compared to TREC DL19.
  • MLDR-zh (Fig. 8): All three methods exhibit a stronger focus on early positions, with fewer blocks selected from later parts. This trend suggests that, for this Chinese dataset, relevant information may be more concentrated in the front positions. However, relevant information are still found and used in other positions, also contributing large portions.
General Trends
Across all datasets, over half of the top 8 selected blocks come from the first 5 positions in documents, reflecting a common positional bias where initial content is often more relevant to the query. Additionally, the figures indicate that relevance distributions vary across datasets. While early positions dominate block selection across datasets, middle and later positions also contribute, indicating that the block selection approaches can retrieve relevant information from throughout the document when necessary.
5.8. Efficiency and Resource Consumption
We compare the reranking time and GPU memory requirements of different models, as shown in Fig. 9. Each experiment is measured on a NVIDIA A100 GPU using the TREC 2019 DL document ranking dataset. As shown in Fig. 9(a), we analyze the reranking time for 100 documents per query (It should be noted the bi-encoder version calculate the block embeddings offline, for fast online search), and in Fig. 9(b), we record the GPU memory usage for various models during training with our settings (see Section 5.3).
(a) Avg. time (seconds) to rerank 100 documets given a query in test set.
(b) GPU memory requirement during training (MB).
Figure 9. Efficiency comparisons on TREC 2019 DL.
5.8.1. Reranking Time
The RankLLaMA model takes the longest time, requiring 14.07 seconds to rerank 100 documents. In contrast, the KeyB2 models were significantly faster. The KeyB2(Llama2)bi model takes 6.25 seconds which is less than half of original time, while the KeyB2(Llama3)cross model require 8.56 seconds. Among all models, the BM25 block matching strategy is the fastest, completing the task in 6 seconds.
5.8.2. GPU Memory Consumption
In terms of GPU memory usage, the RankLLaMA model consumes 33123 MB even with batch size 1, making it the most resource-intensive model. In comparison, The KeyB2 models significantly reduce memory requirements The KeyB2(Llama2)BM25 model used 22523 MB, while the KeyB2(Llama3)cross consume 25297 MB. This represents a substantial memory reduction, ranging from 68% to 76%, demonstrating the efficiency of the KeyB2 models.
These results highlight that KeyB2 models, particularly the KeyB2(Llama2)BM25, offer faster reranking times and lower memory consumption, making them highly efficient and easily deployable for long document retrieval tasks.
5.8.3. Efficiency-Effectiveness Figures
Last, to compare different models’ effectiveness with efficiency together, we draw Fig. 10 to illustrate the overall performance on TREC 2019 DL dataset.
(a) Comparison of speed and effectiveness.
(b) Comparison of memory usage and effectiveness.
Figure 10. Effectiveness with efficiency comparisions on TREC 2019 DL (left and top should be better).
It can be seen the proposed KeyB2 models with different block selection strategies all locate at more left and top parts of the Fig. 10(a) and Fig. 10(b), compared with RankLLaMA model, indicating they have lower reranking latencies, reduced GPU memory requirements, and higher effectiveness than RankLLaMA. If one focuses on the limitation of GPU, one can select to use cross-encoder version. If one focuses on the reranking latency, one can select to use the BM25 or bi-encoder versions, which balance the speed and effectiveness.
6. Conclusion
In this paper, after exploring attention mechanisms in LLMs for IR and providing insights into the inner workings of decoder-only LLMs, we introduce KeyB2, which incorporates block selection strategies with LLMs to improve long document ranking. By selectively focusing on the most relevant blocks, which is similar to the human relevance judgement process, KeyB2 enhances retrieval accuracy while significantly reducing computational overhead.
Our experiments on various datasets, including TREC 2019 DL document ranking task, Robust04, and MLDR-zh (Chinese), demonstrate that KeyB2 models consistently outperform baseline methods like RankLLaMA, BM25 and KeyB. The models achieve the best improvements with different configurations depending on the dataset: Llama3 with cross-encoder showed the highest gains on English datasets, while the newly proposed bi-encoder configuration delivered superior results on MLDR-zh. Furthermore, our analysis of block selection patterns shows a preference for early document positions across datasets, while demonstrating that KeyB2 effectively retrieves relevant content from various positions as needed, ensuring accurate ranking. Additionally, KeyB2 models show consistent effectiveness in zero-shot scenarios, on the unseen Robust04 dataset. After fine-tuning the performances are better.
Overall, KeyB2 achieves better ranking accuracy while processing fewer tokens than approaches that input entire documents, resulting in faster reranking speed and lower GPU memory usage. These gains make KeyB2 a better solution for high-performance IR task, which is one of the key contributions of this paper: reducing the input size and noise while maintaining the LLMs’ effectiveness is possible by intelligently selecting key blocks from documents à la KeyB. Finally, our design of different strategies successfully improves the performances, pointing to new directions in using LLMs for document retrieval.